Compute the weighted average cost of capital (WACC) if the pre-tax cost of debt is 5%, the cost of equity is 10%, the corporate tax rate is 38%, the market value of the firms debt is $120 million, and the market value of the firms equity is %180 million
Answer #1

After-tax cost of debt=5*(1-tax rate)

=5*(1-0.38)=3.1%

Total market value=(120+180)=$300 million

WACC=Respective costs*Respective weight

=(120/300*3.1)+(180/300*10)

which is equal to

=7.24%